# Resource-efficient Inference with Foundation Model Programs

### v1

@@ Abstract
The inference-time resource costs of large language and vision models present a growing challenge in production deployments. We propose the use of foundation model programs, i.e., programs that can invoke foundation models with varying resource costs and performance, as an approach to this problem. Specifically, we present a method that translates a task into a program, then learns a policy for resource allocation that, on each input, selects foundation model “backends” for each program module. The policy uses smaller, cheaper backends to handle simpler subtasks, while allowing more complex subtasks to leverage larger, more capable models. We evaluate the method on two new “streaming” visual question-answering tasks in which a system answers a question on a sequence of inputs, receiving ground-truth feedback after each answer. Compared to monolithic multi-modal models, our implementation achieves up to 98% resource savings with minimal accuracy loss, demonstrating its potential for scalable and resource-efficient multi-modal inference  [^1] Source code and benchmarks are available at https://github.com/Flitternie/FMProgramming..
@@ 1 Introduction
Foundation models (FMs) have reshaped the landscape of machine learning over the past few years, demonstrating unprecedented capabilities in language understanding (Achiam et al., 2023; Dubey et al., 2024), complex reasoning (Lu et al., 2024; Gupta et al., 2024), and multi-modal tasks (Li et al., 2022a; Liu et al., 2024). While much of the community’s attention has focused on their training costs, the inference-time resource use of FMs is increasingly becoming a practical bottleneck. For commercial applications that require real-time responses — for instance, continuous streams of user queries to a multi-modal large language model (MLLM) — computational overhead and high latency can severely degrade user experience and inflate operational expenses Xu et al. (2024).
In this paper, we propose the use of foundation model programs (FMPs) — code in Python-like languages that can call into a variety of specialized vision and language models as subroutines — as a way to address this problem. Such programs have been previously motivated on the basis of the interpretability and flexibility they bring to multi-step tasks (Surís et al., 2023; Gupta & Kembhavi, 2023; Subramanian et al., 2023). Our insight is that they can also enable fine-grained decisions about resource allocation: simpler subtasks can rely on smaller, cheaper backends while more complex components can leverage larger, more capable models.
Concretely, we propose a framework of resource-efficient foundation model programming in which a task is automatically translated into an FMP that captures subtask dependencies and conditional control flow. Each submodule of the program is then assigned to one of several backend models, differing in resource cost and capability. For example, consider the scenario in Figure 1, in which a visual question answering (VQA) system receives the query “Is there a cat sitting or laying on a laptop keyboard?” Here, our method generates a program that uses a small, inexpensive object detection model to check whether both a cat and a laptop are present. Only if that condition is met does it invoke a more powerful vision-language model (VLM) for finer-grained reasoning.
We specifically focus on “streaming” tasks in which the system repeatedly solves a task — for example, answering a question — on a sequence of inputs. Each answer is provided without prior knowledge of the ground truth, and the system receives ground-truth feedback after each answer. In such settings, the cost of using a monolithic model is proportional to the number of inputs processed. By contrast, our approach uses the feedback from the early answers to learn a policy that dynamically selects which backend model to invoke for each subtask, conditioned on the program input. Specifically, we use a combination of a structured REINFORCE estimator and Thompson Sampling to learn this policy.
@@
  • We propose the use of foundation model programs as a flexible approach to cost-efficient inference for complex, multi-modal workflows.
@@ 2 Problem Formulation
[…]
@@ 3 Methodology
Our approach to solving Problem 1 includes an offline code generation phase and an online resource allocation phase.
@@ 3.1 Offline Code Generation
We begin by synthesizing a foundation model program $p$ from the user specification using an LLM. This process produces a task-specific program sketch, including a sequence of generic neural function calls $\{f_{k_{1}},f_{k_{2}},\ldots,f_{k_{N}}\}$, which defines the high-level structure of the computation, while the backend selection for these neural functions is determined online.
@@ 3.2 Online Resource Allocation
This sub-reward decomposes the global reward $R(\vec{v}_{t},x_{t},y_{t})$ into local contributions. It integrates the local computational cost $\mathcal{C}(m_{k_{i},j_{i}})$ associated with backend $m_{k_{i},j_{i}}$, and a portion of the predictive loss $\mathcal{L}(p(x_{t}|\vec{v}_{t}),y_{t})$, that is determined once the entire configuration $\vec{v}_{t}$ is set.
@@
This structured design simplifies the decision space: rather than jointly optimizing over all $n_{k_{1}}\times\cdots\times n_{k_{N}}$ backend combinations, we train $N$ separate subpolicies. Each subpolicy is specialized to one of the $N$ function calls, thereby simplifying the optimization process, enabling parallel learning, and reducing unwanted interference across calls.
@@
     Execute program $p$ with selected configuration $\vec{v}_{t}=(m_{k_{1},j_{1}},...,m_{k_{N},j_{N}})$
@@
The overall framework is detailed in Algorithm 1. Given an input, the structured policy predicts rewards for different backend choices at each function call, incorporating a gradient-based Thompson Sampling to balance exploration and exploitation. This structured framework ensures modular optimization, enabling independent learning of function-specific policies while maintaining tractability. We also provide a no-regret theoretical guarantee for the algorithm in Appendix A.
@@ 4 Benchmark
Motivated by the need for structured, sequential evaluation beyond the single-image-per-query setups typical of existing visual question answering (VQA) datasets Goyal et al. (2017), we introduce two novel streaming VQA benchmarks.
@@ 5.2 Baselines
We evaluate our system against the state-of-the-art multi-modal LLMs (MLLMs) that integrate both vision and language reasoning capabilities Bai et al. (2025).
LLM Routing
As an alternative adaptive strategy, a multi-armed bandit dynamically routes user queries to multi-modal LLMs of varying sizes with different cost-accuracy trade-offs based on estimated rewards, balancing exploration and exploitation Nguyen et al. (2024); Li (2025). However, it does not account for the task structures in user queries.
@@
Following the standard practice in prior works Hu et al. (2024); Jitkrittum et al. (2025), we employ a straightforward yet effective Pareto-random routing strategy through linear interpolation. We implement this approach separately for two scenarios: multi-modal LLMs and static FM program configurations.
@@ 5.3 Results
Compared to the LLM routing baseline, FM Programming achieves a 6% improvement in F1 score at similar computational cost (10% of the largest MLLM). While its maximum F1 score is lower than that of the Single MLLM and LLM routing baselines, this is due to conservative thresholds in the object detection modules that prioritize precision over recall to minimize costly false positives. As shown in Figure 3, the learned policy increasingly favors precision over recall as the cost budget grows. In practice, this trade-off can be adjusted by tuning the reward function used in Structured REINFORCE.
@@
Streaming Open-form VQA Results
@@ 6.1 Multi-modal Reasoning
Multi-modal reasoning tasks, such as visual question answering, require integrating vision and language to address complex queries. End-to-end models, powered by large pre-trained transformers Deitke et al. (2024); Bai et al. (2025), achieve high accuracy but often come with significant computational costs and limited interpretability. In contrast, modular approaches like ViperGPT Surís et al. (2023) and VisProg Gupta & Kembhavi (2023) decompose complex tasks into smaller, executable programs, enhancing flexibility and interpretability by dynamically composing models based on task needs. Compared to the existing methods that rely on fixed programs, our work extends this modular paradigm with a particular focus on cost-efficiency, introducing dynamic FM backend selection to handle varying input complexities for achieving the Pareto frontier between resource and performance.
@@ 6.2 Foundation Model Inference Cost Optimization
Optimizing foundation models, such as LLMs and MLLMs, is essential for deployment in resource-constrained environments. Common techniques include model distillation Hinton et al. (2015); Gu et al. (2023), quantization Gholami et al. (2022); Lin et al. (2024), and speculative decoding Leviathan et al. (2023); Miao et al. (2023). More recently, Compound AI systems have emerged, leveraging multiple model backends to optimize inference costs. Among these, LLM routing selects the best model for a query to balance cost and quality Lu et al. (2023); Hu et al. (2024); Cascading uses a series of models, starting with simpler ones for easy tasks and escalating to complex ones as needed Chen et al. (2023); Nie et al. (2024). Although effective, these methods generally operate at a model-centric level and lack integration with task-specific reasoning structures. In contrast, our approach combines task decomposition with dynamic backend selection in FM programs, enabling resource allocation that adapts to both input complexity and task structures for cost saving.
@@ 7 Conclusion
We introduced the first framework to use foundation model programs in optimizing the trade-off between task performance and resource consumption. By decomposing complex multi-modal reasoning tasks into programs with control flow and dynamically selecting FM backends in online setups, our method can navigate the Pareto frontier of tradeoffs between resource use and performance in real time, in a way that exploits both task and input structure. As demonstrated in our experiments, this strategy can deliver substantial computational savings with minimal performance degradation.
@@ Appendix B Detailed Experimental Setups
All experiments are conducted on a single machine equipped with 8 NVIDIA A40 GPUs (48GB memory each), running CUDA 12.4.
@@
Detailed package versions are listed in the environment file available at https://github.com/Flitternie/FMProgramming/environment.yml. Hyperparameter settings are specified in the experimental configuration files for binary VQA and open-form VQA.
@@ C.2 Streaming Open-form VQA
To assess model robustness and reasoning precision, we incorporate unanswerable images that are visually coherent but semantically invalid with respect to the query. These include both unrelated (random) images and images that are intentionally crafted to closely resemble answerable cases, making them more difficult to distinguish. This setup challenges models not only to infer the correct answer when possible but also to recognize when a question cannot be answered from the image.

### v2

@@ Abstract
The inference-time resource costs of large language and vision models present a growing challenge in the deployment of agentic systems. We propose the use of foundation model programs, i.e., programs that can invoke foundation models with varying resource costs and performance, as an approach to this problem. Specifically, we present a method that translates a task into a program, then learns a policy for resource allocation that, on each input, selects foundation model “backends” for each program module. The policy uses smaller, cheaper backends to handle simpler subtasks, while allowing more complex subtasks to leverage larger, more capable models. We evaluate the method on two new “streaming” visual question-answering tasks in which a system answers a question on a sequence of inputs, receiving ground-truth feedback after each answer. Compared to monolithic multi-modal models, our implementation achieves up to 98% resource savings with minimal accuracy loss, demonstrating its potential for scalable and resource-efficient multi-modal inference  [^1] Source code and benchmarks are available at https://github.com/Flitternie/FMProgramming..
@@ 1 Introduction
Foundation models (FMs) have reshaped the landscape of machine learning over the past few years, demonstrating unprecedented capabilities in language understanding (Achiam et al., 2023; Dubey et al., 2024), complex reasoning (Lu et al., 2024; Gupta et al., 2024), and multimodal tasks (Li et al., 2022a; Liu et al., 2024). While much of the community’s attention has focused on their training costs, the inference-time resource use of FMs is increasingly becoming a practical bottleneck. For commercial applications that require real-time responses — for instance, continuous streams of user queries to a multimodal large language model (MLLM) — computational overhead and high latency can severely degrade user experience and inflate operational expenses Xu et al. (2024).
In this paper, we propose the use of foundation model programs (FMPs) — code in Python-like languages that can call into a variety of specialized vision and language models as subroutines — as a way to address this problem. Such programs have been previously motivated by the interpretability and composability they bring to multi-step tasks (Sur´ıs et al., 2023; Gupta & Kembhavi, 2023; Subramanian et al., 2023) and agentic AI systems (Wu et al., 2023; Khattab et al., 2023). Our insight is that they can also enable fine-grained decisions about resource allocation: simpler subtasks can rely on smaller, cheaper backends while more complex components can leverage larger, more capable models.
Concretely, we propose a framework of resource-efficient foundation model programming in which a task is automatically translated into an FMP that captures subtask dependencies and conditional control flow. The framework is modality-agnostic, making it directly applicable to any domain — text, vision, speech, or multimodal — as long as the task can be expressed as an FMP. Each submodule of the program is then assigned to one of several backend models, differing in resource cost and capability. For example, consider the scenario in Figure 1, in which a visual question answering (VQA) system receives the query “Is there a cat sitting or laying on a laptop keyboard?” Here, our method generates a program that uses a small, inexpensive object detection model to check whether both a cat and a laptop are present. Only if that condition is met does it invoke a more powerful vision-language model (VLM) for finer-grained reasoning.
We specifically focus on “streaming” tasks in which the system repeatedly solves a task — for example, answering a question — on a sequence of inputs. An answer is generated without prior knowledge of the ground truth, then the system receives ground-truth feedback after each answer. In such settings, the cost of using a monolithic model is proportional to the number of inputs processed. By contrast, our approach uses the feedback from the early answers to learn a policy that dynamically selects which backend model to invoke for each subtask, conditioned on the program input. Specifically, we use a combination of a structured REINFORCE estimator and Thompson Sampling to learn this policy.
@@
  • We propose the use of foundation model programs as a flexible approach to cost-efficient inference for complex, multimodal workflows.
@@ 2 Problem Formulation
Now, assume that we analyze the program and produce an arbitrary ordering for the total of $N$ calls to generic neural functions, such that $k_{i}$ is the identity of the generic neural function associated with the $i$-th call in the program. Note that a program can call the same neural function multiple times with different inputs or arguments. The list of neural functions called by the program is then $\langle f_{k_{1}},\;f_{k_{2}},\;\ldots,\;f_{k_{N}}\rangle$. In our example program, the list is $\langle\texttt{ImagePatch.find()},\texttt{ImagePatch.find()},\texttt{ImagePatch.vqa()}\rangle$.
@@ 3 Methodology
Our approach to addressing this problem consists of two key phases: offline code generation and online resource allocation.
@@ 3.1 Offline Code Generation
We begin by synthesizing a foundation model program $p$ from the user specification using an LLM. This process yields a task-specific program sketch, including a sequence of generic neural function calls $\{f_{k_{1}},f_{k_{2}},\ldots,f_{k_{N}}\}$, which defines the high-level structure of the computation, while the backend selection for these neural functions is determined online.
The correctness of the FMP is critical, as it defines the entire computational workflow for the task. To ensure transparency and verifiability, FMPs are constructed with an explicit neurosymbolic representation, making them straightforward to inspect, validate, and debug. Although we employ an LLM to generate these programs, our framework is designed to be agnostic to the source of the FMP and fully supports human-authored workflows.
@@ 3.2 Online Resource Allocation
This sub-reward decomposes the global reward $R(\vec{v}_{t},x_{t},y_{t})$ into structured contributions. It integrates the local computational cost $\mathcal{C}(m_{k_{i},j_{i}})$ associated with backend $m_{k_{i},j_{i}}$, and a portion of the predictive loss $\mathcal{L}(p(x_{t}|\vec{v}_{t}),y_{t})$, that is determined when the program $p$ is executed with the entire configuration $\vec{v}_{t}$.
@@
This structured design simplifies the decision space: rather than jointly optimizing over all $n_{k_{1}}\times\cdots\times n_{k_{N}}$ backend combinations, we train $N$ separate subpolicies. Each subpolicy is specialized to one of the $N$ function calls, thereby simplifying the optimization process, enabling parallel learning, and reducing unwanted interference across calls. This per-call decomposition is critical for a precise credit assignment. In a monolithic joint policy, it is difficult to attribute rewards correctly, especially in the presence of control flow where some function calls may be conditionally skipped. Our structured approach resolves this by updating a subpolicy only when its corresponding function is executed, ensuring that reward signals are correctly aligned with the decisions that caused them.
@@
  Execute program $p$ with selected configuration $\vec{v}_{t}=(m_{k_{1},j_{1}^{*}},...,m_{k_{N},j_{N}^{*}})$
@@
The overall framework is detailed in Algorithm 1 with a walk-through example in Appendix Figure 5. This methodology provides a tractable solution to the online resource allocation challenge, with a theoretical no-regret guarantee provided in Appendix A. By separating offline synthesis from online optimization and employing a decomposed policy structure, we achieve both flexibility and efficiency in program execution.
@@ 4 Benchmark
Motivated by the need for structured, sequential evaluation beyond the single-image-per-query setups typical of existing visual question answering (VQA) datasets Goyal et al. (2017), we introduce two novel streaming VQA benchmarks. These benchmarks are designed to reflect real-world latency-critical deployments by simulating input streams, long-tail difficulty, class imbalance, and adversarial distractors. While partially synthetic, they are constructed using real data (e.g., COCO, GQA) and validated through LLM and human annotations to ensure complexity and robustness.
@@
Figure 2: Examples of the synthetic images in the Streaming Open-form VQA benchmark.
@@ 5.2 Baselines
We evaluate our system against the state-of-the-art multimodal LLMs (MLLMs) that integrate both vision and language reasoning capabilities Bai et al. (2025).
MLLM Routing
As an alternative adaptive strategy, a multi-armed bandit dynamically routes user queries to multimodal LLMs of varying sizes with different cost-accuracy trade-offs based on estimated rewards, balancing exploration and exploitation Nguyen et al. (2024); Li (2025). However, it does not account for the task structures in user queries.
@@
Following the standard practice in prior works Hu et al. (2024); Jitkrittum et al. (2025), we employ a straightforward yet effective Pareto-random routing strategy through linear interpolation. We implement this approach separately for two scenarios: multimodal LLMs and static FM program configurations.
@@ 5.3 Results
Compared to the LLM routing baseline, FM Programming achieves a 6% improvement in F1 score at similar computational cost (10% of the largest MLLM). While its maximum F1 score is lower than that of the Single MLLM and LLM routing baselines, this is due to conservative thresholds in the object detection modules that prioritize precision over recall to minimize costly false positives. As shown in Appendix Figure 8, the learned policy increasingly favors precision over recall as the cost budget grows. In practice, this trade-off can be adjusted by tuning the reward function used in Structured REINFORCE.
@@
Streaming Open-form VQA
@@
Ablation Study
To assess our structured policy design, we compare it against an unstructured policy variant that jointly optimizes all backend selections. As shown in Figures 3 and 4, the unstructured policy (yellow line) significantly underperforms on both benchmarks, due to its inability to manage conditional branches and accurately assign rewards to local decisions, particularly when the resource budgets increase. This highlights the critical role of our Structured REINFORCE algorithm in enabling precise credit assignment and effective resource allocation.
@@ 6.1 Multi-modal Reasoning
Multi-modal reasoning tasks, such as visual question answering, require integrating vision and language to address complex queries. End-to-end models, powered by large pre-trained transformers Deitke et al. (2024); Bai et al. (2025), achieve high accuracy but often come with significant computational costs and limited interpretability. In contrast, modular approaches like ViperGPT Sur´ıs et al. (2023) and VisProg Gupta & Kembhavi (2023) decompose complex tasks into smaller, executable programs, enhancing flexibility and interpretability by dynamically composing models based on task needs. Compared to the existing methods that rely on fixed programs, our work extends this modular paradigm with a particular focus on cost-efficiency, introducing dynamic FM backend selection to handle varying input complexities for achieving the Pareto frontier between resource and performance.
@@ 6.2 Foundation Model Inference Cost Optimization
Optimizing foundation models, such as LLMs and MLLMs, is essential for deployment in resource-constrained environments. Common techniques include distillation Hinton et al. (2015); Gu et al. (2023) and quantization Gholami et al. (2022); Lin et al. (2024) that operate at the model level, and speculative decoding Leviathan et al. (2023); Miao et al. (2023) and mixture-of-experts routing Fedus et al. (2022); Huang et al. (2024) that operate at the token level. In contrast, our approach works at the task level, allowing dynamic resource allocation across multiple subtask backends, making it complementary to these strategies.
More recently, Compound AI systems have emerged, leveraging multiple model backends to optimize inference costs. Among these, LLM routing selects the best model for a query to balance cost and quality Lu et al. (2023); Hu et al. (2024); Cascading uses a series of models, starting with simpler ones for easy tasks and escalating to complex ones as needed Chen et al. (2023); Nie et al. (2024). Although effective, these methods generally lack integration with task-specific reasoning structures. In contrast, our approach combines task decomposition with dynamic backend selection in FM programs, enabling resource allocation that adapts to both input complexity and task structures for cost saving.
@@ 7 Conclusion
We introduced the first framework to use foundation model programs in optimizing the trade-off between task performance and resource consumption. By decomposing complex, multi-modal reasoning tasks into programs with control flow and dynamically selecting FM backends in online setups, our method can navigate the Pareto frontier of trade-offs between resource use and performance in real-time, in a way that exploits both task and input structure. As demonstrated in our experiments, this strategy can deliver substantial computational savings with minimal performance degradation.
@@
Limitations
Our framework’s performance relies on several assumptions: (a) The initial program structure must be correct – an illogical workflow will lead to sub-optimal performance, regardless of backend allocation; (b) The policy is constrained by the capabilities of the available FM backends. If the backend pool is too weak for a given subtask, even optimal routing may not achieve the desired accuracy, as shown in Appendix Figure 9; (c) The online learning algorithm requires immediate ground-truth feedback, which may not be available in real-world settings; (d) We also do not explicitly model error propagation, where an early mistake from a lightweight module can impact downstream steps.
Acknowledgement
We thank the anonymous reviewers for their valuable comments and suggestions in enhancing the paper. This research was supported by the NSF under grant numbers CCF-1918651, 2008240, 2131294, and 2212557, NSF PPoSS award #2316161, ARO award #W911NF-21-1-0009, DARPA award #HR00112320018, NIH CTSA award #UM1TR004906, and US DOT Tier-1 UTC CYBER-CARE grant #69A3552348332.
@@ Appendix B Detailed Experimental Setups
All experiments are conducted on a single machine equipped with 8 NVIDIA A40 GPUs (48GB memory each), running CUDA 12.4. Detailed package versions are listed in the environment file available at GitHub. Hyperparameter settings are specified in the experimental configuration files for binary VQA and open-form VQA.
For the offline code generation, we use a prompt format similar to ViperGPT, including the DSL, the user query, and a one-shot example. In practice, the generated programs are >>90% correct, and errors are easily caught during dry runs or audits. A full prompt example is also available at GitHub.
@@
Figure 5: A walk-through example showing how an FMP is compiled and routed with online resource allocation in Algorithm 1. Here we look at a specific input $x_{t}$ for illustration but the user inputs actually arrive continuously in a stream.
@@ C.2 Streaming Open-form VQA
To assess model robustness and reasoning precision, we incorporate unanswerable images that are visually coherent but semantically invalid with respect to the query. These include both unrelated (random) images and images that are intentionally crafted to closely resemble answerable cases, making them more difficult to distinguish, as illustrated in Figure 2. This setup challenges models not only to infer the correct answer when possible but also to recognize when a question cannot be answered from the image.